OpenStudy (anonymous):

factor completely: see below:

OpenStudy (anonymous):

\[x ^{2}+8x+12\]

OpenStudy (anonymous):

(x+6)(x+2)??

OpenStudy (anonymous):

yup seems good :)

OpenStudy (anonymous):

\[(x+6)(x+2) = x^2 +2x +6x + 12 = x^2 +8x + 12\]

OpenStudy (anonymous):

(x+6)(x+2)

OpenStudy (anonymous):

^ yup that is what @Algebra2013 had :) and was correct.